OpenStudy (anonymous):

Solve using any method. 2x^2 +x-5=0

OpenStudy (anonymous):

Thats supposed to =0 not +0

OpenStudy (jhannybean):

edit your question by the little "edit" button at top

OpenStudy (jhannybean):

But anyways, Group your x terms together with a parenthesis around them\[(2x^2+x)+5=0\]Now from the parenthesis, pull out that leading coefficient, \(2\).\[2\left(x^2+\frac{x}{2}\right) +5\]Remember, if you pull out the 2, you have to keep the equation constant by pulling it out from every variable inside the parenthesis.

OpenStudy (jhannybean):

Are you with me so far?

OpenStudy (anonymous):

so far!

OpenStudy (anonymous):

wait are you meaning to put +5?

OpenStudy (jhannybean):

Good :) it gets a little tricky now. you want to "complete the square" inside the parenthesis, and you have to remember that whatever you do INSIDE the parenthesis, you must do OUTSIDE the parenthesis to keep your equation constant.

OpenStudy (jhannybean):

Sorry, meant to put \(-5\).

OpenStudy (anonymous):

okay so how do i "complete the square"?

OpenStudy (jhannybean):

Let's rewrite that as \[2\left(x^2+\frac{x}{2}\right)-5\]

OpenStudy (jhannybean):

So now "completing the square" means pulling the equation we have inside our parenthesis into quadratic form, \(ax^2+bx+c=0\) So far, we only have \(ax^2+bx\) inside our parenthesis. We need to find our \(c\) value.

OpenStudy (jhannybean):

Still with me? :)

OpenStudy (anonymous):

i think so..

OpenStudy (jhannybean):

\[2\left( \color{red}{x^2+\frac{x}{2}}\right)-5\]\[~~~~~~~~\downarrow\]\[~~~~~\color{red}{ax^2+bx}+c\]

OpenStudy (jhannybean):

Does that kind of clarify it a little bit more?

OpenStudy (anonymous):

yes! thanks!

OpenStudy (jhannybean):

WOOO!!!

OpenStudy (anonymous):

i learn better seeing!(:

OpenStudy (jhannybean):

\[b=\frac{1}{2} \implies c=\left(\frac{\dfrac{1}{2}}{2}\right)^2=\left(\frac{1}{4}\right)^2 =\frac{1}{16}\]

OpenStudy (jhannybean):

Sorry, deleted a little bit of the previous post, \[c=\left(\frac{b}{2}\right)^2\]

OpenStudy (anonymous):

did you get the 2 from the front of the original equation after the parenthesis were added?

OpenStudy (jhannybean):

Nope, follow the post above yours :) \(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[c=\left(\frac{b}{2}\right)^2\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

ohhh!

OpenStudy (jhannybean):

Making sense?

OpenStudy (anonymous):

so far!

OpenStudy (jhannybean):

GREAT! (Trust me, a lot of people have trouble with this method)

OpenStudy (jhannybean):

Now we go back to our original equation, and add in that c value :)\[2\left(x^2+\frac{x}{2} +\color{blue}{\frac{1}{16}}\color{red}{-\frac{1}{16}}\right)-5\color{red}{-\frac{1}{16}}=0\]

OpenStudy (jhannybean):

To stay consistent with our original equation, like I mentioned previously, we must subtract the same value both inside and outside the parenthesis.

OpenStudy (anonymous):

answering my question before i ever ask it!(:

OpenStudy (jhannybean):

OH NO!! MISTAKE!

OpenStudy (jhannybean):

This is something you will have to catch when writing it out,. See how there is a \(2\) factored OUTSIDE the parenthesis?

OpenStudy (anonymous):

yeah?

OpenStudy (jhannybean):

In order to "properly" pull out the \(-\dfrac{1}{16}\), we MUST multiply it to the 2. Just remember, everything that stays inside the parenthesis stays the same, if we pull out anything, then we must take into consideration tnhe number on the outside.

OpenStudy (jhannybean):

Therefore, it would become this: \[2\left(x^2+\frac{x}{2} +\color{blue}{\frac{1}{16}}\color{red}{-\frac{1}{16}}\right)-5\color{red}{-\frac{1}{8}}=0\] and DEFINITELY NOT \[2\left(x^2+\frac{x}{2} +\color{blue}{\frac{1}{16}}\color{red}{-\frac{1}{16}}\right)-5\color{red}{-\frac{1}{16}}=0\]See the difference?

OpenStudy (anonymous):

okay!

OpenStudy (jhannybean):

Now we can simplify what is inside the parenthesis to \[2\left(x^2+\frac{x}{2}+\frac{1}{16}\right) - \frac{81}{16}=0\]\[2\left(x+\frac{1}{4}\right)^2-\frac{81}{16}=0\]

OpenStudy (jhannybean):

And here is a little trick. Remember: \[b=\frac{1}{2} \implies c=\left(\frac{\dfrac{1}{2}}{2}\right)^2=\left(\frac{1}{4}\right)^2 =\frac{1}{16}\] That part?

OpenStudy (jhannybean):

Tell me if I am going too fast btw.

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

No, perfect speed!

OpenStudy (jhannybean):

We can simplify \(\color{blue}{\text{Originally Posted by}}\) @Jhannybean Now we can simplify what is inside the parenthesis to \[2\left(\color{red}{x^2+\frac{x}{2}+\frac{1}{16}}\right) - \frac{81}{16}=0\]\[2\left(\color{red}{x+\frac{1}{4}}\right)^2-\frac{81}{16}=0\] \(\color{blue}{\text{End of Quote}}\) The highlighted portion of that formula by analyzing our \(c\) value :D \[b=\frac{1}{2} \implies c=\left(\frac{\dfrac{1}{2}}{2}\right)^2=\left(\color{red}{\frac{1}{4}}\right)^2 =\frac{1}{16}\] That highlighted red portion of our c value will always be what our quadratic simplifies to

OpenStudy (jhannybean):

It's great because you don't have to think about it!

OpenStudy (anonymous):

oh cool!

OpenStudy (jhannybean):

Is it making sense? :)

OpenStudy (anonymous):

slowly but surely

OpenStudy (jhannybean):

What part are you stuck on? Let's move on from there.

OpenStudy (anonymous):

Not stuck just wondering whats next!

OpenStudy (jhannybean):

Ohh! Ok.

OpenStudy (anonymous):

then are we going to take away the ^2?

OpenStudy (jhannybean):

And there's that typo again -.-

OpenStudy (jhannybean):

It's funny because I'm explaining this to you and i'm making my own mistakes as I go along with this explanation.

OpenStudy (anonymous):

It's okay! Still helping!!

OpenStudy (jhannybean):

\[2\left(x^2+\frac{x}{2} +\color{blue}{\frac{1}{16}}\color{red}{-\frac{1}{16}}\right)-5\color{red}{-\frac{1}{8}}=0\]\[2\left(x^2+\frac{x}{2}+\frac{1}{16}\right)-\frac{41}{8}=0\]\[2\left(x+\frac{1}{4}\right)^2 = \frac{41}{8}\]The reason I was getting \(-\dfrac{81}{16}\) is because I was doing what I told you not to do, and that is pulling out the \(-\dfrac{1}{16}\) from the parenthesis without multiplying it by 2.

OpenStudy (jhannybean):

Now we divide both sides of our equation by \(2\) to isolate the \(\left(x+\dfrac{1}{4}\right)^2\)

OpenStudy (jhannybean):

\[\left(x+\frac{1}{4}\right)^2=\frac{\dfrac{41}{8}}{2}\]\[\left(x+\frac{1}{4}\right)^2 =\frac{41}{16}\] Making sense?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

yes!

OpenStudy (jhannybean):

Just let me know when you're ready to move on.

OpenStudy (jhannybean):

Alright :)

OpenStudy (anonymous):

im ready (:

OpenStudy (jhannybean):

Take the square root of both sides, you end up with: \[x+\frac{1}{4} = \pm \frac{\sqrt{41}}{4}\]\[x= -\frac{1}{4}\pm \frac{\sqrt{41}}{4} = -\frac{1\pm \sqrt{41}}{4}\]

OpenStudy (jhannybean):

And if you'd like to check your answer: http://www.wolframalpha.com/input/?i=2x%5E2%2Bx-5%3D0

OpenStudy (anonymous):

So would that be the final answer?

OpenStudy (jhannybean):

Yep :)

OpenStudy (anonymous):

wow, thanks a lot!!

OpenStudy (jhannybean):

The alternative method is using your equation: \(2x^2+x-5=0\) into the quadratic formula: \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] where \(a = 2~,~ b = 1~,~ c = -5\) You would also end up with \(x= \dfrac{1}{4}(-1\pm \sqrt{41})\) :)